OpenStudy (anonymous):

There are 5 boys and 5 girls on a co-ed basketball team. A team plays 5 players on the court at one time. We can play at most 3 boys. Position doesn’t matter.

OpenStudy (dan815):

okay so consider when you have 0,1,2,3 boys

OpenStudy (dan815):

0 boys = 5 girls = 5C5= 1 1 boy = 4 girls = 5C4 * 5C1=5 2 boys =3 girls = 5C3*5C2= 3 boys =2 girls = 5C2*5C3=

OpenStudy (anonymous):

I think you should first choose 2 girls, then chose 3 players from both girls and boys, somehow.

OpenStudy (anonymous):

It is apparently okay for the team to have no boys.

OpenStudy (anonymous):

I wonder if \({5\choose 2}{3+5\choose 3}\) is equal to what dan's method.

OpenStudy (zarkon):

they are not the same...you would be double counting

OpenStudy (anonymous):

\[ {10 \choose 5} - {5 \choose 5} - {5 \choose 4}{5 \choose 1} \]Hmm, I got to figure out where my double counting is coming from.

OpenStudy (anonymous):

that is just counting by complement. The double counting came from the fact that the set of 5 girls and the set of 8 people remaining are not disjoint

OpenStudy (anonymous):

Hmm, so the same girl can be chosen twice?

OpenStudy (anonymous):

yes. (5C2) (8C3) would count at least 1 girl (twice?) maybe more, but counting some girls again is certain